how many grams of sulfur are in 5.53g sample of al2(so3)3
1
Expert's answer
2020-10-04T09:18:07-0400

Molar mass of "Al_2(SO_3)_3=294 \ g."

So , "294\ g" of "Al_2(SO_3)_3" Contain "96\ g" of sulfur

"1\ g" of "Al_2(SO_3)_3" will have "\frac{96}{294}\ g" of sulfur.

"5.53\ g" of "Al_2(SO_3)_3" will have "\frac{96}{294}\times 5.53= 1.80\ g" of sulfur